calculate where f(x)=g(x)from your graph and hence solve the inequality


1
Expert's answer
2021-09-01T06:03:01-0400

This question is incomplete but we can take an example as per the question asked.

Let "f(x)=x^2, g(x)=x"

On plotting these on the graph, we get



From the graph, "f(x)=g(x)" will obtained where these two curves intersect each other.

So, "f(x)=g(x)" at "(0,0)" and "(1,1)" i.e., at "x=0, y=0" and at "x=1, y=1" .

Now, for "f(x)>g(x)"

"\Rightarrow x^2>x
\\\Rightarrow x^2-x>0
\\\Rightarrow x(x-1)>0
\\x\in(-\infty,0)\cup(1,\infty)"

Now, for "f(x)<g(x)"

"\Rightarrow x^2<x
\\\Rightarrow x^2-x<0
\\\Rightarrow x(x-1)<0
\\x\in(0,1)"